OpenStudy (anonymous):

Use implicit differentiation to find the derivative of x(1+y)^(1/2) + y(1+2x)^(1/2) = 2x

OpenStudy (anonymous):

\[x \sqrt{1+y}+y \sqrt{1+2x} = 2x\]

OpenStudy (carniel):

o.o very hard..

OpenStudy (anonymous):

Sorry.

zepdrix (zepdrix):

Oh you appear to be offline now :) If you come back on, lemme know if you still need help with this.